three times the greater of two consecutive odd integers is twice the least by 37
Answer by mananth(16949)

let the numbers be x & x+2
3 times greater = 3(x+2)
twice the least = 2x
...
3(x+2)=2x+37 ( i have presumed it is +37)
3x+6=2x+37
3x-2x=37-6
x= 31
..
the numbers are 31 & 33
...
m.ananth@hotmail.ca